A person is standing with their feet spread 1.2m apart. Their legs are 1 m long, and their weight is equally divided between their legs. What minimal static coefficient of friction is required so that they do not have to use their muscle to prevent their legs from sliding outward (so they don't do a split)?
Answer #1

The weight is divided between the legs

F = =Mu*mg = mg*1/1.2

Mu = 1/1.2 = 0.83

Static coefficient of friction = 0.83